A delivery truck is transporting boxes of two sizes: large and small. The combined weight of a large box and a small box is 70 pounds. The truck is transporting 65 large boxes and 60 small boxes. If the truck is carrying a total of 4425 pounds in boxes, how much does each type of box weigh?
Answer by ankor@dixie-net.com(22740)

write equations for each statement
A delivery truck is transporting boxes of two sizes: large and small.
let b = the no. of large boxes
let s = the no. of small boxes
:
The combined weight of a large box and a small box is 70 pounds.
b + s = 70
we can use this form for substitution
s = (70-b)
:
The truck is transporting 65 large boxes and 60 small boxes.
If the truck is carrying a total of 4425 pounds in boxes
65b + 60s = 4425
Simplify, divide by 5
13b + 12s = 885
:
how much does each type of box weigh?
Replace s with (70-b) in the above equation
13b + 12(70-b) = 885
13b + 840 - 12b = 885
13b - 12b = 885 - 840
b = 45 large boxes
then
70 - 45 = 25 small boxes
;
:
Check:
65*45 = 2925
60*25 = 1500
---------------
total wt:4425